OpenStudy (anonymous):

how do you divide numbers into decimals (without a calculator)? e.g. 15 divided by 6

OpenStudy (anonymous):

Well, 6 fits into 15 two full times. So the answer will be 2 point something. We have 3 leftover from teh 15, so that's 2 and 3/6 3/6 is half, so we get 2.5

OpenStudy (anonymous):

:O thank you!

OpenStudy (anonymous):

Think of highest common factors 15=5*3 6=3*2 15/6=5/2 =2.5

OpenStudy (anonymous):

So what about a question like 18 divided by 30?

OpenStudy (anonymous):

You will just have to do simplification. 18/30 = 9/15 = 3/5 = 0.6

OpenStudy (anonymous):

18=6*3 30=6*5 18/30=3/5=6/10=0.6

OpenStudy (amistre64):

its called long hand division: ? ---- 6 | 15 ; what is the number 6(n) that is closest to, but is not greater than, 15? 6(1) = 6 6(2) = 12 6(3) = 18 <- to big, go with '2' 2. ? ---- 6|15.0000 -12 <- subtract 6(2) to get a new result to work on ------ 30 <- bring down the next number, in this case '0' now the strategy is to do the same step, try to find a number, 6(n), that is close to but less than 30. 6(1)=6 6(2)=12 6(3)=18 6(4)=24 6(5)=30 <- an exact math, use '5' 2.5 ---- 6|15.0000 -12 ------ 30 -30 <- subtract 6(5) to get a new result ---- 0 <- our new number is z zero which means we have reached the end. 2.5 is our answer.

OpenStudy (amistre64):

i spose 'less than or equal to' is better advice lol